# Extending Stable and Popular Matching Algorithms from Bipartite to Arbitrary Instances

### v1

@@ 3.1 Maximum Size $\gamma$-Stable Fractional-Matching
Let $V=\{v_{1},\dots,v_{n}\}$ be an indexing of the vertices. We describe the fine-tuned first step of the meta-algorithm for this problem in more detail.
For each edge $e=(v_{i},v_{j})\in E$, $i<j$ we create four parallel edges $e_{1},e_{2},e_{3},e_{4}$. For $v_{i}$, $e_{1}$ is considered the best copy of $e$, while $e_{2}$ is second best, $e_{3}$ is third and $e_{4}$ is last. For $v_{j}$, $e_{4}$ is the best copy of $e$, $e_{3}$ is second, $e_{2}$ is third and $e_{1}$ is last.
We create the strict preferences of a vertex $v$ as follows. For any edges $e,f$, the worst copy of $f$ (which may be either $f_{1}$ or $f_{4}$) is always ranked worse than any best, second or third copy of $e$. Furthermore, the second best copy of $f$ is better than the best copy of $e$ if and only if $p_{v}(f)\geq p_{v}(e)+\gamma_{f}^{v}$ and the third copy of $f$ is better than the best copy of $e$ if and only if $p_{v}(f)\geq p_{v}(e)+\delta_{f}^{v}$. Within the best and last copies, the edges are ranked according to a strict extension of $v$’s original weak preferences. This can be obtained in the following way. For a vertex $v\in V$ and $e=(u,v)$ let $a(e)$ be the best, $b(e)$ be the second, $c(e)$ be the third and $d(e)$ be the last copy. Set $p_{v}(a(e))=p_{v}(e)$, $p_{v}(b(e))=p_{v}(e)-\gamma_{e}^{v}$ and $p_{v}(c(e))=p_{v}(e)-\delta_{e}^{v}$. Then, for each vertex, these new (possibly negative) preference numbers define a weak order, in which we break ties in a way such that among edges with the same value the $c$ copies are best, the $b$ copies are second and the $a$ copies are last. Finally, the worst copies are ranked strictly last in the order they are in $p_{v}()$, breaking the ties arbitrarily.
@@
Proof: Suppose for contradiction that there is a $\gamma$-blocking edge $e=(v_{i},v_{j})$ to $M$. Suppose by symmetry that $p_{v_{i}}(e)\geq p_{v_{i}}(M)+\gamma_{e}^{v_{i}}$ and $p_{v_{j}}(e)\geq p_{v_{j}}(M)+\delta_{e}^{v_{j}}$. Take the second best copy of $e$ for $v_{i}$, which is the third best copy of $e$ for $v_{j}$. This copy of $e$ blocks $M^{\prime}$ by the definition of the rankings, contradiction.
$\blacksquare$
@@
For any $\gamma$-stable half-matching $N$ it holds that $\sum_{e\in E}N(e)\leq\frac{3}{2}\sum_{e\in E}M(e)$.
@@
Each stable half-matching $N$ defines a matching $N^{\prime\prime}$ in $G^{\prime\prime}$ as follows. (By a result of Tan[33], we know that this consists of disjoint value $1$ edges and cycles of value $\frac{1}{2}$ edges). If $N(e)=1$ for $e=(v_{i},v_{j})$, then we add $(v_{i},v_{j}^{\prime})$ and $(v_{i}^{\prime},v_{j})$ to $N^{\prime\prime}$. If we have a cycle of value $\frac{1}{2}$ edges on vertices $\{v_{i_{1}},\dots,v_{i_{k}}\}$ with $i_{1}<\cdots<i_{k}$, then we add $(v_{i_{1}},v_{i_{2}}^{\prime}),(v_{i_{2}},v_{i_{3}}^{\prime}),\dots,(v_{i_{k}},v_{i_{1}}^{\prime})$.
@@
In the first case, we get that there is an edge $(v_{i},v_{j})$ such that $v_{i}$ and $v_{j}$ are both unsaturated in $M$, contradicting the $\gamma$-stability of $M$. In the second case, let the component be $\{(v_{i_{1}},v_{i_{2}}^{\prime}),(v_{i_{2}}^{\prime},v_{i_{3}}),(v_{i_{3}},v_{i_{4}}^{\prime})\}$, where $(v_{i_{2}}^{\prime},v_{i_{3}})$ is the edge of $M^{\prime\prime}$. We have that $i_{1},i_{2},i_{3}$ are distinct and so are $i_{2},i_{3},i_{4}$, but it may happen that $i_{1}=i_{4}$.
Since $v_{i_{1}}$ and $v_{i_{4}}^{\prime}$ are unmatched by $M^{\prime\prime}$, $v_{i_{1}}$ and $v_{i_{4}}$ are unsaturated in $M^{\prime}$. Since the best copy of $(v_{i_{1}},v_{i_{2}})$ for $v_{i_{2}}$ does not block $M^{\prime}$, $v_{i_{2}}$ is saturated with first, second or third copies. Since the best copy of $(v_{i_{3}},v_{i_{4}})$ for $v_{i_{3}}$ does not block $M^{\prime}$, $v_{i_{3}}$ is saturated with first, second or third copies. Hence, either the second or the third best copy of $e=(v_{i_{2}},v_{i_{3}})$ for $v_{i_{2}}$ has positive weight in $M^{\prime}$. Suppose by symmetry it is the second best copy. Then, for the same reason, we must have $p_{v_{i_{2}}}(e)\geq p_{v_{i_{2}}}(v_{i_{1}},v_{i_{2}})+\gamma_{e}^{v_{i_{2}}}$ and $p_{v_{i_{3}}}(e)\geq p_{v_{i_{3}}}(v_{i_{3}},v_{i_{4}})+\delta_{e}^{v_{i_{3}}}$. This contradicts the $\gamma$-stability of $N$. $\blacksquare$
We also show that if the instance has no odd preference cycles, then the algorithm outputs a weakly stable matching that approximates the maximum weakly stable matching. Recall that this means that in every odd cycle, there is an agent that prefers its clockwise neighbour to its other neighbour, and there is another agent that prefers its anti-clockwise neighbour to its other neighbour.
@@
If the instance does not admit odd preference cycles, then the algorithm outputs an integral weakly stable matching $M$, such that $|M|\geq\frac{2}{3}|N|$ for any other weakly stable matching $N$.
@@
By the result of Tan on stable partitions [33], we know that if there is no odd preference cycles, then each weakly stable half-matching consists of disjoint value 1 edges and even cycles of value $\frac{1}{2}$ edges. Also, in this case, Tan’s algorithm always outputs an integral weakly stable matching. Hence, our algorithm also outputs an integral matching.
Finally, we have already shown that $|M|\geq\frac{2}{3}|N|$ in Theorem 1, so this also clearly holds for weakly stable integral matchings too. ∎
By Theorems 1 and 2 we have shown that both variants of finding a maximum size weakly stable fractional-matching mentioned in the introduction has a $\frac{3}{2}$-approximation algorithm.
@@ 3.2 Maximum Size Popular Fractional-Matching
In this section we solve the maximum size popular fractional-matching problem, Max-PRI for short. Recall that for this problem, we assumed that we have strict preference orders $\succ_{v}$ for each agent $v\in V$.
We describe the fine-tuned first step of the meta-algorithm. For each edge $e=(v_{i},v_{j})$, $i<j$, we create two parallel edges $e_{1},e_{2}$. Agent $v_{i}$ considers $e_{1}$ the good copy and $e_{2}$ the bad one, while $v_{j}$ considers $e_{2}$ the good and $e_{1}$ the bad one. The new strict preferences are created such that for any vertex $v_{i}$, it first ranks the good copies in their original order and strictly after then the bad copies in their original order.
@@
Max-PRI is solvable in polynomial time. Furthermore, the output of the algorithm is half-integral, extra-popular and has maximum size even among the barely-defendable fractional-matchings.
@@
We decompose $\phi$ as follows. First, we define $\phi^{0}$ by setting $\phi_{v}^{0}(e,e)=0$ and $\phi_{v}^{0}(e,f)=\phi_{v}(e,f)$ for $e\neq f$. We also set $M^{0}(e)=M(e)-\phi_{v}(e,e)=M(e)-\phi_{u}(e,e)$ and $N^{0}(e)=N(e)-\phi_{v}(e,e)$ for all $e\in E$. We can do this, as $\mathrm{vote}_{v}(e,e)$ is always 0, so this does not change $\Delta(M,N,\phi)$, i.e. $\Delta(M,N,\phi)=\Delta(M^{0},N^{0},\phi^{0})$. We also must still have that for each $e=(v_{i},v_{j})$, $\sum\limits_{f\in E(v_{i})\cup\{\emptyset\}}\phi^{0}_{v_{i}}(e,f)=\sum\limits_{f\in E(v_{j})\cup\{\emptyset\}}\phi^{0}_{v_{j}}(e,f)=M^{0}(e)$ and $\sum\limits_{f\in E(v_{i})\cup\{\emptyset\}}\phi^{0}_{v_{i}}(f,e)=\sum\limits_{f\in E(v_{j})\cup\{\emptyset\}}\phi^{0}_{v_{j}}(f,e)=N^{0}(e)$, because of the axiom $\phi_{v_{i}}(e,e)=\phi_{v_{j}}(e,e)$ on sensible-pairings.
@@
Let $k=0$. Create a graph $H^{k}=(W^{k},E^{k})$, where each $e=(v_{i},v_{j})\in E$ with $N^{k}(e)>0$ receives two vertices $e_{N}^{i},e_{N}^{j}$ and each with $M^{k}(e)>0$ receives vertices $e_{M}^{i},e_{M}^{j}$. We also add a vertex $v^{\emptyset}$ corresponding to $\emptyset$. Then, we define the edges of $H^{k}$ such that $(e_{M}^{i},f_{N}^{j})$ is an edge if and only if $i=j$ and $\phi^{k}_{v_{i}}(e,f)>0$. We also add edges edge $(e_{M}^{i},e_{M}^{j})$ and $(e_{N}^{i},e_{N}^{j})$ for each $e=(v_{i},v_{j})$. Finally, we add an edge $(e_{M}^{i},v^{\emptyset})$ if and only if $\phi^{k}_{v_{i}}(e,\emptyset)>0$ and an edge $(e_{N}^{i},v^{\emptyset})$ if $\phi_{v_{i}}(\emptyset,e)>0$. Let $F^{k}=\{(e_{M}^{i},e_{M}^{j})\mid e\in E,\;M^{k}(e)>0\}\cup\{(e_{N}^{i},e_{N}^{j})\mid e\in E,\;N^{k}(e)>0$}.
@@
It is clear that either there is an $F^{k}-E^{k}\setminus F^{k}$ alternating cycle in $H^{k}$ or an almost $F^{k}-E^{k}\setminus F^{k}$ alternating cycle, which alternates at each vertex except $v^{\emptyset}$. This holds because each vertex apart from $v^{\emptyset}$ is adjacent to at least one edge from both $F^{k}$ and $E^{k}\setminus F^{k}$, as we have (and will maintain) for each $e=(v_{i},v_{j})$ that $\sum\limits_{f\in E(v_{i})\cup\{\emptyset\}}\phi^{k}_{v_{i}}(e,f)=\sum\limits_{f\in E(v_{j})\cup\{\emptyset\}}\phi^{k}_{v_{j}}(e,f)=M^{k}(e)$ and $\sum\limits_{f\in E(v_{i})\cup\{\emptyset\}}\phi^{k}_{v_{i}}(f,e)=\sum\limits_{f\in E(v_{j})\cup\{\emptyset\}}\phi^{k}_{v_{j}}(f,e)=N^{k}(e)$.

### v3

@@ 3.1 Maximum Size $\gamma$-Stable Fractional-Matching
Let $V=\{v_{1},\dots,v_{n}\}$ be an indexing of the vertices.
The fine-tuned first step.
For each edge $e=(v_{i},v_{j})\in E$, $i<j$ we create four parallel edges $e^{1},e^{2},e^{3},e^{4}$. For $v_{i}$, $e^{1}$ is considered the best copy of $e$, while $e^{2}$ is second best, $e^{3}$ is third and $e^{4}$ is last. For $v_{j}$, $e^{4}$ is the best copy of $e$, $e^{3}$ is second, $e^{2}$ is third and $e^{1}$ is last.
We create the strict preferences of a vertex $v$ as follows. For any edges $e,f$, the worst copy of $f$ (which may be either $f^{1}$ or $f^{4}$) is always ranked worse than any best, second or third copy of $e$. Furthermore, the second best copy of $f$ is better than the best copy of $e$ if and only if $p_{v}(f)\geq p_{v}(e)+\gamma_{f}^{v}$ and the third copy of $f$ is better than the best copy of $e$ if and only if $p_{v}(f)\geq p_{v}(e)+\delta_{f}^{v}$. Within the best and last copies, the edges are ranked according to a strict extension of $v$’s original weak preferences. This can be obtained in the following way. For a vertex $v\in V$ and $e=(u,v)$ let $a(e)$ be the best, $b(e)$ be the second, $c(e)$ be the third and $d(e)$ be the last copy. Set $p_{v}(a(e))=p_{v}(e)$, $p_{v}(b(e))=p_{v}(e)-\gamma_{e}^{v}$ and $p_{v}(c(e))=p_{v}(e)-\delta_{e}^{v}$. Then, for each vertex, these new (possibly negative) preference numbers define a weak order, in which we break ties in a way such that among edges with the same value the $c$ copies are best, the $b$ copies are second and the $a$ copies are last. Finally, the worst copies are ranked strictly last in the order they are in $p_{v}()$, breaking the ties arbitrarily.
@@
Proof: Suppose for contradiction that there is a $\gamma$-blocking edge $e=(v_{i},v_{j})$ to $M$. Suppose by symmetry that $p_{v_{i}}(e)\geq p_{v_{i}}(M)+\gamma_{e}^{v_{i}}$ and $p_{v_{j}}(e)\geq p_{v_{j}}(M)+\delta_{e}^{v_{j}}$. Take the second best copy of $e$ for $v_{i}$, which is the third best copy of $e$ for $v_{j}$. This copy of $e$ blocks $M^{\prime}$ by the definition of the rankings, contradiction.  $\blacksquare$
@@
For any $\gamma$-stable half-matching $N$, it holds that $\sum_{e\in E}N(e)\leq\frac{3}{2}\sum_{e\in E}M(e)$.
@@
Each half-matching $N$ defines a matching $N^{\prime\prime}$ in $G^{\prime\prime}$ as follows. If $N(e)=1$ for $e=(v_{i},v_{j})$, then we add $(v_{i},v_{j}^{\prime})$ and $(v_{i}^{\prime},v_{j})$ to $N^{\prime\prime}$. If we have a cycle of value $\frac{1}{2}$ edges on vertices $\{v_{i_{1}},\dots,v_{i_{k}}\}$, then we add $(v_{i_{1}},v_{i_{2}}^{\prime}),(v_{i_{2}},v_{i_{3}}^{\prime}),\dots,(v_{i_{k}},v_{i_{1}}^{\prime})$. Finally, if we have a path of value $\frac{1}{2}$ edges on vertices $\{v_{i_{1}},\dots,v_{i_{k}}\}$, then we add $(v_{i_{1}},v_{i_{2}}^{\prime}),(v_{i_{2}},v_{i_{3}}^{\prime}),\dots,(v_{i_{k-1}},v_{i_{k}}^{\prime})$.
@@
In the first case, we get that there is an edge $(v_{i},v_{j})$ such that $v_{i}$ and $v_{j}$ are both unsaturated in $M$, contradicting the $\gamma$-stability of $M$. In the second case, let the component be $\{e^{\prime\prime}=(v_{i_{1}},v_{i_{2}}^{\prime}),f^{\prime\prime}=(v_{i_{2}}^{\prime},v_{i_{3}}),g^{\prime\prime}=(v_{i_{3}},v_{i_{4}}^{\prime})\}$, where $f^{\prime\prime}=(v_{i_{2}}^{\prime},v_{i_{3}})$ is the edge of $M^{\prime\prime}$. We have that $i_{1},i_{2},i_{3}$ are distinct and so are $i_{2},i_{3},i_{4}$, but it may happen that $i_{1}=i_{4}$.
Since $v_{i_{1}}$ and $v_{i_{4}}^{\prime}$ are unmatched by $M^{\prime\prime}$, $v_{i_{1}}$ and $v_{i_{4}}$ are unsaturated in $M^{\prime}$. Since the best copy of $e$ for $v_{i_{2}}$ does not block $M^{\prime}$, $v_{i_{2}}$ is saturated with first, second or third copies. Since the best copy of $g$ for $v_{i_{3}}$ does not block $M^{\prime}$, $v_{i_{3}}$ is saturated with first, second or third copies. Hence, either the second or the third best copy of $f=(v_{i_{2}},v_{i_{3}})$ for $v_{i_{2}}$ has positive weight in $M^{\prime}$. Suppose by symmetry it is the second best copy. Then, for the same reason, we must have $p_{v_{i_{2}}}(e)\geq p_{v_{i_{2}}}(f)+\gamma_{e}^{v_{i_{2}}}$ and $p_{v_{i_{3}}}(e)\geq p_{v_{i_{3}}}(g)+\delta_{e}^{v_{i_{3}}}$. This contradicts the $\gamma$-stability of $N$.  $\blacksquare$
3.1.1 Maximum Size Weakly Stable Fractional-Matching
For completeness, we describe an even simpler algorithm for the special case Max-fSRTI, i.e. the case, when there are no $\gamma$ numbers, each agent $v$ just has a preference valuation $p_{v}()$, inducing a weak preference order $\succeq_{v}$ ($e\succ_{v}f\Leftrightarrow p_{v}(e)>p_{v}(f)$ and $e\succeq_{v}f\Leftrightarrow p_{v}(e)\geq p_{v}(f)$) and our goal is to find a maximum size weakly stable half-matching.
Then, in the first step of the algorithm, we do the following. For each $e=(v_{i},v_{j}),i<j$, we create three parallel edges $e^{i},e^{0},e^{j}$. For $v_{i}$, $e^{i}$ is best, $e^{0}$ is middle and $e^{j}$ is worst, and for $v_{j}$, $e^{j}$ is best, $e^{0}$ is middle and $e^{i}$ is worst. Then, the preference orders for $v_{i}\in V$ are created from $\succeq_{v_{i}}$ such that for each $e\in E(v_{i})$, $e^{i}\succ e^{0}$, and $f^{0}\succ e^{i}$ if and only if $f$ is strictly preferred to $e$. Within the same copy types, the edges are ranked according to an arbitrary strict extension of $\succeq_{v_{i}}$. The copies $e^{j}$, $j\neq i$ are ranked last in the order $\succeq_{v}$, breaking the ties in an arbitrary way. An easy way to create these preferences is to go through $\succeq_{v_{i}}$, and in each tie $e_{1},..,e_{k}]$ just substitute it with $e^{i}_{1}\succ\cdots e_{k}^{i}\succ e_{1}^{0}\succ\cdots e_{k}^{0}$, then append to worst copies of the edges strictly after them by breaking the ties arbitrarily.
To give an example, let $\succeq_{v_{i}}=e\succ[f,g]\succ h$. Then, we get $e^{i}\succ e^{0}\succ f^{i}\succ g^{i}\succ f^{0}\succ g^{0}\succ h^{i}\succ h^{0}\succ e^{j_{1}}\succ f^{j_{2}}\succ g^{j_{3}}\succ h^{j_{4}}$, where $e=(v_{i},v_{j_{1}}),f=(v_{i},v_{j_{2}}),g=(v_{i},v_{j_{3}}),h=(v_{i},v_{j_{4}})$.
@@ 3.1.1 Maximum Size Weakly Stable Fractional-Matching
Max-fSRTI can be $\frac{3}{2}$-approximated in polynomial time even with the simpler algorithm.
@@
Let $M$ be the output of the simpler algorithm and $M^{\prime}$ be the pre-image of $M$ in $I^{\prime}$.
Claim 3.
The output half-matching $M$ is weakly stable.
Proof: Suppose for contradiction that there is a blocking edge $e=(v_{i},v_{j})$ to $M$. Then, the copy $e^{0}$ blocks $M^{\prime}$ by the definition of the rankings, contradiction.  $\blacksquare$
Claim 4.
For any weakly stable half-matching $N$ it holds that $\sum_{e\in E}N(e)\leq\frac{3}{2}\sum_{e\in E}M(e)$.
Proof: Let $N$ be a maximum size weakly stable half-matching.
Create a bipartite graph $G^{\prime\prime}=(V,V^{\prime},E^{\prime\prime})$, and matchings $N^{\prime\prime},M^{\prime\prime}$ as in Theorem 1.
In the first case, we get that there is an edge $(v_{i},v_{j})$ such that $v_{i}$ and $v_{j}$ are both unsaturated in $M$, contradicting the weak stability of $M$. In the second case, let the component be $\{e^{\prime\prime}=(v_{i_{1}},v_{i_{2}}^{\prime}),f^{\prime\prime}=(v_{i_{2}}^{\prime},v_{i_{3}}),g^{\prime\prime}=(v_{i_{3}},v_{i_{4}}^{\prime})\}$, where $f^{\prime\prime}=(v_{i_{2}}^{\prime},v_{i_{3}})$ is the edge of $M^{\prime\prime}$, i.e. $M(f)>0$. We have that $i_{1},i_{2},i_{3}$ are distinct and so are $i_{2},i_{3},i_{4}$, but it may happen that $i_{1}=i_{4}$.
Since $v_{i_{1}}$ and $v_{i_{4}}^{\prime}$ are unmatched by $M^{\prime\prime}$, $v_{i_{1}}$ and $v_{i_{4}}$ are unsaturated in $M^{\prime}$. Since $e^{i_{2}}$ does not block $M^{\prime}$, only $f^{i_{2}},f^{0}$ can have positive value in $M^{\prime}$. Since $g^{i_{3}}$ does not block $M^{\prime}$, only $f^{i_{3}},f^{0}$ can have positive value in $M^{\prime}$. Hence, we have that $M^{\prime}(f^{0})>0$. Then, since $e^{i_{2}}$ and $g^{i_{3}}$ does not block $M^{\prime}$, we must have $f\succ_{v_{i_{2}}}e$ and $f\succ_{v_{i_{3}}}g$, contradicting the stability of $N$.  $\blacksquare$
@@ 3.2 Maximum Size Popular Fractional-Matching
In this section we solve the maximum size popular fractional-matching problem, Max-fPRI for short. Recall that for this problem, we assumed that we have strict preference orders $\succ_{v}$ for each agent $v\in V$.
The fine-tuned first step.
We describe the fine-tuned first step of the meta-algorithm. For each edge $e=(v_{i},v_{j})$, $i<j$, we create two parallel edges $e_{a},e_{b}$. Agent $v_{i}$ considers $e_{a}$ the good copy and $e_{b}$ the bad one, while $v_{j}$ considers $e_{b}$ the good and $e_{a}$ the bad one. The new strict preferences are created such that for any vertex $v_{i}$, it first ranks the good copies in their original order and strictly after then the bad copies in their original order.
@@
Max-fPRI is solvable in polynomial time. Furthermore, the output of the algorithm is half-integral, extra-popular and has maximum size even among the barely-defendable fractional-matchings.
@@
We decompose $\phi$ as follows. First, we define $\phi^{0}$ by setting $\phi_{v}^{0}(e,e)=0$ and $\phi_{v}^{0}(e,f)=\phi_{v}(e,f)$ for $e\neq f$. We also set $M^{0}(e)=M(e)-\phi_{v}(e,e)=M(e)-\phi_{u}(e,e)$ and $N^{0}(e)=N(e)-\phi_{v}(e,e)$ for all $e\in E$. We can do this, as $\mathrm{vote}_{v}(e,e)$ is always 0, so this does not change $\Delta(M,N,\phi)$, i.e. $\Delta(M,N,\phi)=\Delta(M^{0},N^{0},\phi^{0})$. We also must still have that for each $e=(v_{i},v_{j})$,
  1. 1.
    $\sum\limits_{f\in E(v_{i})\cup\{\emptyset\}}\phi^{0}_{v_{i}}(e,f)=\sum\limits_{f\in E(v_{j})\cup\{\emptyset\}}\phi^{0}_{v_{j}}(e,f)=M^{0}(e)$ and
  2. 2.
    $\sum\limits_{f\in E(v_{i})\cup\{\emptyset\}}\phi^{0}_{v_{i}}(f,e)=\sum\limits_{f\in E(v_{j})\cup\{\emptyset\}}\phi^{0}_{v_{j}}(f,e)=N^{0}(e)$,
because of the axiom $\phi_{v_{i}}(e,e)=\phi_{v_{j}}(e,e)$ on sensible-pairings.
@@
Let $k=0$. Create a graph $H^{k}=(W^{k},E^{k})$, where each $e=(v_{i},v_{j})\in E$ with $N^{k}(e)>0$ receives two vertices $e_{N}^{i},e_{N}^{j}$ and each with $M^{k}(e)>0$ receives vertices $e_{M}^{i},e_{M}^{j}$. We also add a vertex $v^{\emptyset}$ corresponding to $\emptyset$. Then, we define the edges of $H^{k}$ such that $(e_{M}^{i},f_{N}^{j})$ is an edge if and only if $i=j$ and $\phi^{k}_{v_{i}}(e,f)>0$. We also add edges $(e_{M}^{i},e_{M}^{j})$ and $(e_{N}^{i},e_{N}^{j})$ for each $e=(v_{i},v_{j})$. Finally, we add an edge $(e_{M}^{i},v^{\emptyset})$ if and only if $\phi^{k}_{v_{i}}(e,\emptyset)>0$ and an edge $(e_{N}^{i},v^{\emptyset})$ if $\phi_{v_{i}}(\emptyset,e)>0$. Let $F^{k}=\{(e_{M}^{i},e_{M}^{j})\mid e\in E,\;M^{k}(e)>0\}\cup\{(e_{N}^{i},e_{N}^{j})\mid e\in E,\;N^{k}(e)>0$}.
To proceed with the proof, we need the following claim.
Claim 5.
Let $H=(W,E\cup F)$ ($E\cap F=\emptyset$) be a multigraph with $|E\cup F|\geq 1$, where $F$ is an almost perfect matching that covers all but one vertex $v_{0}\in W$ and every vertex $v\in W\setminus\{v_{0}\}$ has degree at least one in $E$. Then, it holds that either there is an EFE-F-alternating closed walk in $H$, or an almost EFE-F-alternating closed walk, which alternates at every vertex except $v_{0}$.
Proof: We prove this by induction on $|W|$. Clearly, as $F$ forms an almost perfect matching in $H$, where only $v_{0}$ is unmatched, we have that $|W|=2m+1$ for some $m\in\mathbb{Z}$. If $|W|=1$, then $W=\{v_{0}\}$, so by $|E\cup F|\geq 1$, there is a loop edge $v_{0}v_{0}$ and the statement follows.
Suppose we know the statement for $|W|=2m+1$ and take the case, where $|W|=2m+3\geq 3$. Then, $F\neq\emptyset$. Let $(u,v)\in F$. If there is an edge $(u,v)\in E$ too, then we are done, since we found a closed EFE-F-alternating cycle. Otherwise, as $u$ and $v$ have degree at least one in $E$, we can create nonempty sets $U\subseteq W$ and $V\subseteq W$, such that each vertex in $U$ is connected to $u$ in $E$ and each vertex in $V$ is connected to $v$ in $E$.
Take the graph $H^{\prime}$, which is the restriction of $H$ to $W\setminus\{u,v\}$, but for each $u^{\prime}\in U^{\prime}$ and $v^{\prime}\in V^{\prime}$, we add an edge $(u^{\prime},v^{\prime})$. We call these edges virtual edges. Let $F^{\prime}$ be the restriction of $F$ to $H^{\prime}$ and $E^{\prime}$ be the other edges. For this graph, we have that $|W^{\prime}|=2m+1$, $F^{\prime}$ forms an almost perfect matching, and every $v\in W^{\prime}\setminus\{v_{0}\}$ has degree at least one in $E^{\prime}$ (if a vertex $w$ was only adjacent to $u$ or $v$, then we created at least one new edge incident to $w$). Since $E^{\prime}\neq\emptyset$, as $|V^{\prime}|\geq 1,|U^{\prime}|\geq 1$, we have by induction that there is an $E^{\prime}-F^{\prime}$-alternating closed walk in $H^{\prime}$ or an almost $E^{\prime}-F^{\prime}$-alternating closed walks, which alternates except from $v_{0}$. From this, we can create an EFE-F-alternating closed walk or an almost EFE-F-alternating closed walk in $H^{\prime}$ by substituting each virtual edge $(u^{\prime},v^{\prime})$ with $(u^{\prime},u),(u,v),(v,v^{\prime}($.  $\blacksquare$
@@
Observe that $F^{k}$ is an almost perfect matching in $H^{k}$ covering all vertices but $v^{\emptyset}$ and each vertex apart from $v^{\emptyset}$ is adjacent to at least one edge from $E^{k}\setminus F^{k}$. The latter holds as we have (and will maintain) properties 1 and 2, meaning that for each $e=(v_{i},v_{j})$ that $\sum\limits_{f\in E(v_{i})\cup\{\emptyset\}}\phi^{k}_{v_{i}}(e,f)=\sum\limits_{f\in E(v_{j})\cup\{\emptyset\}}\phi^{k}_{v_{j}}(e,f)=M^{k}(e)$ and $\sum\limits_{f\in E(v_{i})\cup\{\emptyset\}}\phi^{k}_{v_{i}}(f,e)=\sum\limits_{f\in E(v_{j})\cup\{\emptyset\}}\phi^{k}_{v_{j}}(f,e)=N^{k}(e)$. Hence, using Claim 5 for $H^{k}=(W^{k},E^{k}\setminus F^{k}\cup F^{k})$ and $v_{0}=v^{\emptyset}$ we get that, if $\phi^{k}$ is not the all zero function yet (so $E^{k}\neq\emptyset$ by properties 1, 2 and the construction of $H^{k}$), then either there is an $F^{k}-E^{k}\setminus F^{k}$ alternating closed walk in $H^{k}$ or an almost $F^{k}-E^{k}\setminus F^{k}$ alternating closed walk, which alternates at every vertex except $v^{\emptyset}$.